Calculate the magnitude and direction of the electric field 0.45 m from a +7.85 x 10–9 C point charge.






1
Expert's answer
2022-02-28T10:05:16-0500

Answer


Electric field is given by

"E=\frac{KQ}{r^2}\\=\frac{9*10^9**7.85*10^{-9}}{(0.45)^2}\\=348.89N/m"